| u =u. in a non-magnetic material, the frequency is 100 MHz and the complex internal impedance ñ = 200e 10.5

as expressed.

(a) Calculate the damping constant .

(b) Calculate the phase constant.
Answer #1

* Given, page-1 Ñ = 200 Jong On = 0.59 we know, on- ; tante 0.5= 2 tant (CE) Ž= 2 tant (wo) tan(o)= we We = 0.017455064 when10 9 6 page 2 5 = x 100x10° X 0.017455064 18 we know, I to=8.85% 1012 2 Elm 1836 78 X 107 36T Elm 108x159 xo.017455064 18 J 0Рge - 3 чу у х = 48-48 ( xxis 10-2 - 6 a = 08:486 3x (D хJ2 IDI -S X = 08-03 63 х ( px 12.XT а — Ч8. 1863 x 2 X 2 2 xlos 1 «В = 2 xe Page 4 3 В = 2 x 22 }x 3 В= 2 - 0952 38096 ad/m 211 hadam В= 2Т) 3